OpenStudy (anonymous):

Find the zeros of: f(x) = 2x^2 + 2x + 7

OpenStudy (anonymous):

set the function equal to zero and solve for x

OpenStudy (anonymous):

The answer my teacher gave: \[\frac{-1\pm \sqrt{13}i}{2}\] though i don't think it's right. If it is, can you show me the steps to achieve this?

OpenStudy (anonymous):

2x^2 + 2x + 7 = 0 use quadratic formula to solve for x

OpenStudy (anonymous):

Is the answer I provided correct?

OpenStudy (anonymous):

That seems about right to me, but I think it could have gone further.

OpenStudy (anonymous):

ur answer is correct

OpenStudy (anonymous):

I just don't understand how to get that answer.

OpenStudy (anonymous):

I know to set a,b,c. then solve it out, but I think i'm doing something wrong.

OpenStudy (anonymous):

lol. Sorry, I always factor out the 1/2 But, give me one second I'll show ya!

OpenStudy (anonymous):

Thanks! I'll gladly wait :)

OpenStudy (anonymous):

\[ax^2+bx+c\] quadratic formula: \[x= \frac{ -b+-\sqrt{(b)^2-4(a)(c)} }{ 2a }\] Using the quadratic formula, solve for x. \[x= \frac{ -2+-\sqrt{(2)^2-4(2)(7)} }{ 2*2 }\] \[x= \frac{ -2+-\sqrt{(2)-56} }{ 4 }\] \[x= \frac{ -2+-\sqrt{-52} }{ 4 }\] Now, you have to express \[\sqrt{-52}\] in terms of i \[\sqrt{-52}\]= \[\sqrt{-1}\sqrt{52}\] =\[i \sqrt{52}\] \[x= \frac{ -2+-i\sqrt{52} }{ 4 }\] simplify radicals \[\sqrt{52}=\sqrt{4*13}=\sqrt{2^2*13}=2\sqrt{13}\] \[x= \frac{ -2+-2i\sqrt{13} }{ 4 }\] and the +- is (plus or minus)

OpenStudy (anonymous):

There's more...

OpenStudy (anonymous):

You could factor 2 from \[-2+2i \sqrt{13}\] giving \[2(i \sqrt{13}-1)\] and then doing the same thing on the other one so, \[2(-i \sqrt{13}-1)\] Now, at this time you divide by the 4, and you'll get \[x=\frac{ -(i \sqrt{13)}-1 }{ 2 } or x=\frac{ i \sqrt{13}-1 }{ 2 }\]

OpenStudy (anonymous):

You could also factor out 1/2, but that'd be pointless.

OpenStudy (anonymous):

AND You could have also used completing the square method, which is eight times faster and easier.

OpenStudy (anonymous):

holy crap thank you! I UNDERSTAND!